OpenStudy (anonymous):

solve the following equation for y. 27x+3y=6 I'm helping my sister understand this but it's been so long since i took it! am i doing this right 27x+3y=6 3y=6-27x after that i cant recall the rest

OpenStudy (anonymous):

How do you get y by itself? Divide by 3! \[\frac{3y}{3} = \frac{6 - 27x}{3}\]\[y = 2 - 9x\]

OpenStudy (anonymous):

gotcha, for some reason i was trying to figure out where to leave x at lol thank you!